OpenStudy (anonymous):

Find the volume of the solid formed by rotating the region inside the first quadrant enclosed by y=x^3,y=16x about the x-axis.

OpenStudy (solomonzelman):

your boundaries are \(y=x^3\) and \(y=16x\) and it is rotated about the x,axis correct?

OpenStudy (solomonzelman):

x-axis *

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

that will look like a washer method

OpenStudy (solomonzelman):

do you know the washer method?

OpenStudy (solomonzelman):

(don't be ahsamed not to know, if you don't know it is ok. the only person i can not help is the silent one)

OpenStudy (anonymous):

i do but have trouble with the set up

OpenStudy (solomonzelman):

ok, lets figure the limits of integration together I will name them differently for our convinience \(f(x)=16x\) \(g(x)=x^3\)

OpenStudy (solomonzelman):

lets find the points where these two functions intersect

OpenStudy (solomonzelman):

16x=x³ 16=x² (and disregard negative x solutions because we are only in the 1st quadrant)

OpenStudy (solomonzelman):

x=?

OpenStudy (anonymous):

4

OpenStudy (solomonzelman):

yes, and one more?

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So our limits of integration (for the region bounded by 16x and x³) is (x=0 and x=4) \ (i.e. the intersection points after which the region stops)

OpenStudy (solomonzelman):

So far we know \(\large\color{slate}{\displaystyle\int\limits_{0}^{4}}\)

OpenStudy (anonymous):

okay got it and hw to determine inner vs outer for the formula

OpenStudy (solomonzelman):

yes, and if you were to graph the two functions \(f(x)=16x\) and \(g(x)=x^3\) then for all values on the interval [0,4] (which is the interval we need) it is true that f(x)≥g(x)

OpenStudy (solomonzelman):

So, g(x) is always below the f(x) in our case. right?

OpenStudy (solomonzelman):

(you can test it using any value between 0 and 4)

OpenStudy (anonymous):

okay great thanks , makes more sense now , you were a big help!!

OpenStudy (solomonzelman):

you are thinking me already, we aren;t done yet, aren we?

OpenStudy (solomonzelman):

thanking*